OpenStudy (anonymous):

Steven was given the function, f(x)=1.025^7t and asked to determine the rate of change and the type of exponential function. What is the best way for Steven to describe the function? A. The rate of change is 0.025% and represents exponential decay. B. The rate of change is 1.025% and represents exponential growth. C. The rate of change is 2.5% and represents exponential growth. D. The rate of change is -.025% and represents exponential decay.

OpenStudy (anonymous):

Someone help!

OpenStudy (misty1212):

Hi!!

OpenStudy (anonymous):

Hi! Could you help??

OpenStudy (misty1212):

sure

OpenStudy (misty1212):

\[0.025\] is what number as a percent?

OpenStudy (misty1212):

hint: move the decimal two places to the right !

OpenStudy (anonymous):

2.5?

OpenStudy (misty1212):

no not \(2.5?\) but \(2.5!!!\)

OpenStudy (anonymous):

actually 2.5%

OpenStudy (misty1212):

so \(2.5\%\)is the rate of increase, and of course it is growth because \(1.025>1\)

OpenStudy (misty1212):

and you are done \(\color\magenta\heartsuit\)

OpenStudy (anonymous):

So then it would be C?

OpenStudy (misty1212):

always C!

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

could you help with another one possibly?

OpenStudy (anonymous):

The fee for the parking garage at the airport is $16 for the first day plus $4.50 for each additional day. Which equation represents the fee, f, for leaving your car at the airport for d days, where d is the number of days beyond the first day? A. d = 16f + 4.5 B. f = 4.5d + 16 C. f = 16d + 4.5 D. d = 4.5d + 16